OpenStudy (anonymous):

(4/3)p +2q - 8 = 0 10p + 9q = 3(2p + q) + 24 When I did this: 10p + 9q = 3p + 3q + 24 4p + 6q = 24 4p + 6((2/3)p + 4) = 24 4p + 4p - 24 = 24 8p = 48 p = 6 (4/3)(6) + 2q - 8 = 0 8 + 2q = 8 2q = 0 q = 0 Did I do this right??? Would this be a) (6,6) b) (-6,-6) c) always true d) none of these

OpenStudy (anonymous):

well right off the back i can see you distributed your 3 incorrectly on 10p + 9q = 3(2p + q) + 24 you said this equaled 10p + 9q = 3p + 3q + 24 it is actually 10p + 9q = 6p + 3q + 24

OpenStudy (anonymous):

Oh....XD

OpenStudy (anonymous):

I actually meant 6p. That was just a typo.

OpenStudy (anonymous):

oh i see, looks like you're on the right track then

OpenStudy (anonymous):

I did this correctly then?